Please help me:
Solve and check all proposed solutions.
((12)/(x^2-9)) + ((x-5)/(x-3)) = 0
I am coming up with an answer of x+1/x+3 and it just doesn't seem right. Please help me with this.

Answer by stanbon(75887)

Solve and check all proposed solutions.
((12)/(x^2-9)) + ((x-5)/(x-3)) = 0
------
Multiply thru by x^2-9 to get:
12 + (x-5)(x+3) = 0
---
x^2 -2x -3 = 0
Factor:
(x-3)(x+1) = 0
Possible solutions:
x = 3 or x = -1
----
x=3 is not a solution because it makes the denominator zero.
---
Only solution:
x = -1
=======================
Cheers,
Stan H.